# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
After the movie was over it suddenly came to us that Cherry and Marcia didn’t have a way to get home. Two-Bit gallantly offered to walk them home—the west side of town was only about twenty miles away—but they wanted to call their parents and have them come and get them. Two-Bit finally talked them into letting us drive them home in his car. I think they were still half-scared of us. They were getting over it, though, as we walked to Two-Bit’s house to pick up the car. It seemed funny to me that Socs—if these girls were any example—were just like us. They liked the Beatles and thought Elvis Presley was out, and we thought the Beatles were rank and that Elvis was tuff, but that seemed the only difference to me. Of course greasy girls would have acted a lot tougher, but there was a basic sameness. I thought maybe it was money that separated us.
From S. E. Hinton, The Outsiders. Copyright 1967 by S. E. Hinton
Based on the passage, which of the following is most likely true about Ponyboy?
Choices: [(a) Ponyboy doesn’t want to take the girls home, but he feels obligated to assist them., (b) Cherry is attempting to develop a deep connection with Ponyboy, but he is resistant., (c) Ponyboy wishes he had more money so he could be more like the Socs., (d) This is Ponyboy’s first time having an authentic conversation with a member of the Socs.]
Answer index: 3
Subject: Language Arts
Skill: use-the-correct-verb-with-compound-subjects
Description: Complete the sentence with the best verb.
Whenever Skip runs out into the yard, the chickens and the rooster ___ into the coop.
Choices: [(a) disappears, (b) disappear]
Answer index: 1
Subject: Language Arts
Skill: distinguish-facts-from-opinions
Description: Which sentence states a fact?
@@
Choices: [(a) The spinach plant is native to Asia., (b) Spinach tastes worse than any other vegetable.]
Answer index: 0
Subject: Language Arts
Skill: choose-the-best-transition
Description: Choose the transition that compares, or shows similarities.
@@
Betty helps her community by raising money for a youth center. ___, Dale makes his neighborhood more beautiful by planting trees.
Choices: [(a) Similarly, (b) For example]
Answer index: 0
@@
Subject: Language Arts
Skill: analyze-passages-from-narrative-of-the-life-of-frederick-douglass-part-1
Description: Read the following passage from "Narrative of the Life of Frederick Douglass". In this excerpt, Douglass describes some features found on the plantation where he was enslaved as a young child.
@@
Colonel Lloyd kept a large and finely cultivated garden, which afforded almost constant employment for four men, besides the chief gardener, Mr. M’Durmond. This garden was probably the greatest attraction of the place. During the summer months, people came from far and near—from Baltimore, Easton, and Annapolis—to see it. It abounded in fruits of almost every description, from the hardy apple of the north to the delicate orange of the south. This garden was not the least source of trouble on the plantation. Its excellent fruit was quite a temptation to the hungry swarms of boys, as well as the older slaves, belonging to the colonel, few of whom had the virtue or the vice to resist it.
From Frederick Douglass, Narrative of the Life of Frederick Douglass
What is this passage mainly about?
Choices: [(a) It is about why Douglass enjoyed spending time in the garden and stable on Colonel Lloyd’s plantation., (b) It is about how Colonel Lloyd’s fancy garden and fine horses created challenges for enslaved people., (c) It is about how Colonel Lloyd paid more attention to his garden than to the horses in his stable., (d) It is about why enslaved people preferred to work in Colonel Lloyd’s stable instead of his garden.]
Answer index: 1
Subject: Language Arts
Skill: analyze-short-stories
Description: Read the story.
title: Chiarelli’s Bakery
Ms. Chiarelli nodded impatiently. "Come on! If we hurry, we can make cassatelle too."
Based on the **fourth paragraph**, what is Aram’s attitude about his future career?
Choices: [(a) He is confident that he will achieve success., (b) He is sure that he will fail, but he wants to try., (c) He is uncertain how things will turn out.]
Answer index: 0
Subject: Language Arts
Skill: determine-the-main-idea-of-a-passage
Description: Read the passage.
@@
All over the planet, ancient tombs and temples lie hidden from view. So how do archaeologists know where to begin digging? Thanks to satellites, that question is becoming easier to answer. In the pioneering field of satellite archaeology, satellites four hundred miles above Earth can identify structures buried under soil, covered by modern construction, or obscured by forests. For example, mud bricks used by ancient Egyptians are more dense than soil, and the differences are detected by the satellites. Advanced computer programs map buried settlements and distinguish them from water, vegetation, and farmland. Sarah Parcak, a pioneer in the field, says of the technology, "It allows you to literally strip away vegetation and see entire cities beneath the rain forest canopy. This is the unbelievable future of archaeology."
What is the main, or central, idea of the passage?
Choices: [(a) Archaeologists must determine where ancient structures are buried before they can begin excavating them., (b) Archaeologists are uncovering ancient hidden structures by using satellite imagery.]
Answer index: 1
Subject: Language Arts
Skill: use-parallel-structure
Description: Complete the sentence.
@@
In his novel Nineteen Eighty-Four, George Orwell describes a dystopian London in the hands of a totalitarian government that is paranoid, watchful, and ___.
Choices: [(a) mercilessly oppressive, (b) mercilessly oppresses its citizens]
Answer index: 0
Subject: Language Arts
Skill: use-dictionary-entries-to-determine-correct-usage
Description: Look at this dictionary entry.
dictionary word: exasperate
dictionary pos: verb
dictionary inflections: exasperating, exasperated
dictionary entry: 1 to make very angry; to irritate Marie was exasperated by Barbara’s constant complaining.
dictionary etymology: from Latin (exasperare)
Which sentence uses the word **exasperate** correctly?
Choices: [(a) Mrs. Hampton realized she was exasperating a bad situation when she failed to pay the parking ticket promptly., (b) Gary’s reluctance to do basic household chores exasperated Khalil and Chase.]
Answer index: 1
Subject: Language Arts
Skill: trace-an-argument-set-1
Description: Read the text.
title: The Connected City
@@
The bottom line is that online connectivity is a necessity in today’s high-tech society. Cities across the country are scrambling to keep up with ever-increasing demands for Wi-Fi by the general public. In fact, an Internet Society survey of more than ten thousand Internet users around the world found that eighty-three percent of the respondents agreed that access to the Internet should be considered a basic human right. Whether viewed as a public service or as a way to boost the local economy, people expect advanced technology on the go. Travelers, business owners, and local residents should be able to rely on the local government to help them stay connected.
What is the author’s main claim or argument?
Choices: [(a) Access to free Wi-Fi throughout a city helps to promote tourism., (b) Free Wi-Fi in cities benefits residents as well as the local economy., (c) Only low-income residents of urban areas should be provided with free Wi-Fi.]
Answer index: 1
Subject: Language Arts
Skill: which-text-is-most-formal
Description: Which expression of thanks is more formal?
@@
Choices: [(a) We’re so pleased you could join us on our special day. Thank you again for the wonderful gift., (b) We’re super happy that you could be with us on our special day. Thanks again for the awesome gift!]
Answer index: 0
Subject: Language Arts
Skill: trace-an-argument-set-2
Description: Read the text.
title: Fact or Fiction?
@@
Looking beyond the author’s own life events, memoir can inform readers about the world in the same way that other nonfiction can. For example, a memoir written by the head chef of a famous restaurant is likely to educate the reader about the inner workings of a professional kitchen. A memoir composed by a cello player in an esteemed orchestra will probably help the reader learn more about music. And a memoir recalling an adolescence during the 1980s may describe pop culture during that time. Memoir has a way of relaying facts about anything from an occupation to brief fashion trends, all of it meaningful to the author.
What is the author’s main claim or argument?
Choices: [(a) Memoir is best categorized under the umbrella of literary nonfiction., (b) Memory is too unreliable for memoir to be considered a type of nonfiction., (c) While many memoirs are fact-based, many authors are untrustworthy., (d) The most profitable memoirs are those shown to be the most factual.]
Answer index: 0
Subject: Language Arts
Skill: choose-the-topic-sentence-that-best-captures-the-main-idea
Description: Review the details below.
Cell phones can be used to cheat on tests by sending answers to friends or by accessing the Internet.
Texting during class is like passing notes and is a waste of time.
Choose the best topic sentence to introduce a paragraph containing these details.
@@
Choices: [(a) Cell phones in the classroom can have a negative impact on student learning., (b) There are many legitimate arguments for allowing students to have cell phones in class, even if they are disruptive., (c) Cell phones have become a near constant presence at school and at home.]
Answer index: 0
Subject: Language Arts
Skill: identify-sentence-fragments
Description: Is there a sentence fragment?
@@
Before signing in, you must first register for an account. Unless you’re already a member of the site.
Choices: [(a) yes, (b) no]
Answer index: 0
@@
Subject: Language Arts
Skill: use-context-as-a-clue-to-the-meanings-of-foreign-expressions
Description: What is the meaning of the foreign expression in bold?
@@
In 2014, giant moths swarmed **en masse** in parts of Southeast Asia. More than eight hundred sightings were reported in Singapore, while thousands disrupted a soccer match at Darul Makmur Stadium in Malaysia.
Choices: [(a) coincidentally, (b) as a group, (c) unexpectedly, (d) fatally]
Answer index: 1
Subject: Language Arts
Skill: use-the-correct-foreign-expression
Description: Complete the text with the correct foreign expression.
@@
The Norwood Bistro features an eclectic ___ menu so that diners can select a variety of dishes.
Choices: [(a) à la carte, (b) crème de la crème]
Answer index: 0
Subject: Language Arts
Skill: identify-audience-and-purpose
Description: Review the fund-raising pamphlet and then answer the question.
@@
Be a hero in an animal’s life! Donate to the Martin County Animal Shelter. Adoption fees cover less than eight percent of our operating expenses. We rely on donations to ensure that every animal receives excellent care. Your generous gift provides food, shelter, and veterinary care to homeless and rescued pets. Please donate today!
Which best indicates that the primary audience is "community members who care about animals"?
Choices: [(a) The writer indicates that donating is a way for readers to "be a hero in an animal’s life.", (b) The writer argues that the shelter is necessary to keep stray animals off the streets and prevent illnesses., (c) The writer asks readers to donate time as well as money., (d) The writer includes information about funding sources and operating expenses.]
Answer index: 0
Subject: Language Arts
Skill: classify-figures-of-speech-euphemism-hyperbole-oxymoron-paradox
Description: Which figure of speech is used in this text?
This movie is so predictable that I’m likely to **die of boredom** if I keep watching it.
Choices: [(a) paradox, (b) hyperbole]
Answer index: 1
Subject: Language Arts
Skill: analyze-short-stories-set-1
Description: Read the story.
title: Lake Magic
Gramps appeared with dry towels. "I’ve made popcorn and a warm fire. And your mom called. She’s fine." Gramps paused, then added, "Everybody’s fine."
What is the main reason Izzie is looking for Champ?
Choices: [(a) She hopes Champ will take her brother’s place as a friend to her., (b) She wants to prove to her grandfather that she can find the legendary lake monster., (c) She wants Champ’s "lake magic" to make her family the way it used to be.]
Answer index: 2
Subject: Language Arts
Skill: explore-words-with-new-or-contested-usages
Description: Which text uses the word **nauseous** in its traditional sense?
@@
Choices: [(a) Everyone seemed to enjoy the magnolia-scented candle, but it made Walter feel rather **nauseous**., (b) Everyone seemed to enjoy the magnolia-scented candle, but Walter found the smell rather **nauseous**.]
Answer index: 1
Subject: Language Arts
Skill: analyze-the-development-of-informational-passages-set-2
Description: Read the text.
title: Manzanar: A Site of Conscience
Adapted from the National Park Service, "Manzanar: A Site of Conscience"
What is the main focus of the text?
@@
Choices: [(a) how the attack on Pearl Harbor led to war against Japan, (b) the history of the internment of Japanese Americans, with a focus on Manzanar, (c) how Japanese Americans adjusted to everyday life at Manzanar, (d) the history of injustices against Japanese Americans, including internment at Manzanar]
Answer index: 1
Subject: Language Arts
Skill: analyze-the-development-of-informational-passages-set-1
Description: Read the text.
title: Language Change
Adapted from the National Science Foundation, "Language and Linguistics: Language Change"
What is the main focus of the text?
Choices: [(a) why the works of Shakespeare and Chaucer are difficult for modern readers, (b) why different languages evolve at different rates, (c) how and why language changes over time]
Answer index: 2
Subject: Language Arts
Skill: choose-punctuation-to-avoid-fragments-and-run-ons
Description: Which is the best way to complete the text?
@@
Zoe traced her family’s lineage back to the eighteenth ___ cross-referenced her grandmother’s stories with a genealogical index at the library.
Choices: [(a) century. She, (b) century, furthermore, she, (c) century, she]
Answer index: 0
Subject: Language Arts
Skill: transition-logically-between-claims-evidence-analysis-and-counterclaims
Description: Complete the text with the transition that best connects **the two pieces of supporting evidence**.
@@
Parents should limit how much screen time their children are allowed to have each day. A study found that children whose parents limited screen time experienced physical benefits, including better sleep habits and decreased risk of obesity. ___, these children earned better grades in school, were more social and helpful, and were less aggressive with peers.
Choices: [(a) In other words, (b) Moreover, (c) For example]
Answer index: 1
Subject: Language Arts
Skill: choose-the-analysis-that-logically-connects-the-evidence-to-the-claim
Description: Read the claim and the supporting evidence.
*Claim:** Hunter is remarkably virtuous.
*Evidence:** Hunter is generous with his time.
Why does the evidence support the claim? Choose the **analysis** that better explains the connection.
Choices: [(a) Generosity is a key virtue., (b) Hunter believes generosity need not involve money.]
Answer index: 0
Subject: Language Arts
Skill: identify-run-on-sentences
Description: Is this a run-on sentence?
@@
Clinging to a two-point lead in the final minute of the fourth quarter, the Tigers triumphed when their defense recovered a fumble and sealed the victory.
Choices: [(a) yes, (b) no]
Answer index: 1
Subject: Language Arts
Skill: use-words-accurately-and-precisely
Description: Complete the text with the better phrase.
Karen and Steven are ___ on a documentary film project about the effects of drought on local wildlife.
Choices: [(a) collaborating together, (b) working together]
Answer index: 1
Subject: Language Arts
Skill: classify-figures-of-speech-review
Description: Which figure of speech is used in this text?
He had hoped to find his missing watch, but **he couldn’t find the time**.
Choices: [(a) onomatopoeia, (b) pun]
Answer index: 1
Subject: Language Arts
Skill: avoid-double-illogical-and-unclear-comparisons
Description: Which text best completes the sentence?
As a dancer, I studied tap and ballet. In my experience, tap was ___.
Choices: [(a) easier to learn than any form of dance, (b) easier to learn, (c) easiest to learn]
Answer index: 1
Subject: Language Arts
Skill: choose-the-best-evidence-to-support-a-claim
Description: Consider this claim:
Online colleges prepare students for careers just as well as traditional brick-and-mortar schools do.
Which is the strongest evidence to support the claim?
@@
Choices: [(a) "I’ve hired two people recently who received their college degrees through online programs, and the quality of their performance is as good as, if not better than, that of their peers," reported business owner Fred Browning., (b) Online college programs often provide career preparation services that are equivalent to those provided by traditional college programs, including interview preparation, job fairs, and even internship placement., (c) Many students choose to gain some credits through an online college program and then successfully transfer to a brick-and-mortar institution to complete a degree.]
Answer index: 1
@@
Subject: Language Arts
Skill: analyze-short-stories-set-2
Description: Read the story.
title: Water Rising
Polara looked to the horizon, at the unmistakable sparkle of the approaching sea. She wasn’t sure how brave she felt, but she was willing. "I am," she answered.
Based on the **first and second paragraphs**, what internal conflict does Polara struggle with?
Choices: [(a) the duty to stand watch versus her inclination to daydream, (b) the pull of the unfamiliar versus her desire to stay at Shublik, (c) respect for Commander Axel versus her irritation at his order, (d) the enjoyment of the view versus her fear about the water on the horizon]
Answer index: 1
Subject: Language Arts
Skill: use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing
Description: Which statement primarily appeals to **logos**, or reason, to support the position that **more money does not always lead to increased happiness**?
@@
Choices: [(a) Like most people, I enjoy the feeling of buying a new gadget, but it turns out that the happiness that we get from a new purchase doesn’t last long., (b) While it’s true that a lack of money can cause stress and unhappiness, it does not necessarily follow that money will lead to more happiness after basic needs are met.]
Answer index: 1
Subject: Language Arts
Skill: choose-the-word-whose-connotation-and-denotation-best-match-the-sentence
Description: Complete the sentence with the word that best fits the overall meaning and tone.
Dan appreciated the gallery owner’s ___ efforts to include artwork from the local community.
Choices: [(a) calculating, (b) deliberate]
Answer index: 1
Subject: Language Arts
Skill: evaluate-counterclaims
Description: Consider this claim:
Veggie burgers are every bit as tasty and satisfying as regular beef burgers.
Now consider how someone might argue against this. Which sentence presents the stronger and more reasonable counterclaim to the above claim?
@@
Choices: [(a) Veggie burgers have no taste whatsoever and cannot compare in any way to a delicious beef burger., (b) Regular beef burgers have more good-tasting fat and filling protein than veggie burgers.]
Answer index: 1
Subject: Language Arts
Skill: use-etymologies-to-determine-the-meanings-of-words

### v2

@@ Appendix D Summary of Skills
Subject: Language Arts Skill: analyze-passages-from-the-outsiders-part-1 Description: Read the following passage from "The Outsiders". In this excerpt, Ponyboy, the narrator, and his friend, another greaser named Two-Bit, are walking with two girls from a rival group called the Socs. After the movie was over it suddenly came to us that Cherry and Marcia didn’t have a way to get home. Two-Bit gallantly offered to walk them home—the west side of town was only about twenty miles away—but they wanted to call their parents and have them come and get them. Two-Bit finally talked them into letting us drive them home in his car. I think they were still half-scared of us. They were getting over it, though, as we walked to Two-Bit’s house to pick up the car. It seemed funny to me that Socs—if these girls were any example—were just like us. They liked the Beatles and thought Elvis Presley was out, and we thought the Beatles were rank and that Elvis was tuff, but that seemed the only difference to me. Of course greasy girls would have acted a lot tougher, but there was a basic sameness. I thought maybe it was money that separated us. From S. E. Hinton, The Outsiders. Copyright 1967 by S. E. Hinton Based on the passage, which of the following is most likely true about Ponyboy? Choices: [(a) Ponyboy doesn’t want to take the girls home, but he feels obligated to assist them., (b) Cherry is attempting to develop a deep connection with Ponyboy, but he is resistant., (c) Ponyboy wishes he had more money so he could be more like the Socs., (d) This is Ponyboy’s first time having an authentic conversation with a member of the Socs.] Answer index: 3
Subject: Language Arts Skill: use-the-correct-verb-with-compound-subjects Description: Complete the sentence with the best verb. Whenever Skip runs out into the yard, the chickens and the rooster ___ into the coop. Choices: [(a) disappears, (b) disappear] Answer index: 1
@@
Subject: Language Arts Skill: distinguish-facts-from-opinions Description: Which sentence states a fact? Choices: [(a) The spinach plant is native to Asia., (b) Spinach tastes worse than any other vegetable.] Answer index: 0
@@
Subject: Language Arts Skill: choose-the-best-transition Description: Choose the transition that compares, or shows similarities. Betty helps her community by raising money for a youth center. ___, Dale makes his neighborhood more beautiful by planting trees. Choices: [(a) Similarly, (b) For example] Answer index: 0
@@
Subject: Language Arts Skill: analyze-passages-from-narrative-of-the-life-of-frederick-douglass-part-1 Description: Read the following passage from "Narrative of the Life of Frederick Douglass". In this excerpt, Douglass describes some features found on the plantation where he was enslaved as a young child. Colonel Lloyd kept a large and finely cultivated garden, which afforded almost constant employment for four men, besides the chief gardener, Mr. M’Durmond. This garden was probably the greatest attraction of the place. During the summer months, people came from far and near—from Baltimore, Easton, and Annapolis—to see it. It abounded in fruits of almost every description, from the hardy apple of the north to the delicate orange of the south. This garden was not the least source of trouble on the plantation. Its excellent fruit was quite a temptation to the hungry swarms of boys, as well as the older slaves, belonging to the colonel, few of whom had the virtue or the vice to resist it. From Frederick Douglass, Narrative of the Life of Frederick Douglass What is this passage mainly about? Choices: [(a) It is about why Douglass enjoyed spending time in the garden and stable on Colonel Lloyd’s plantation., (b) It is about how Colonel Lloyd’s fancy garden and fine horses created challenges for enslaved people., (c) It is about how Colonel Lloyd paid more attention to his garden than to the horses in his stable., (d) It is about why enslaved people preferred to work in Colonel Lloyd’s stable instead of his garden.] Answer index: 1
Subject: Language Arts Skill: analyze-short-stories Description: Read the story. title: Chiarelli’s Bakery Ms. Chiarelli nodded impatiently. "Come on! If we hurry, we can make cassatelle too." Based on the **fourth paragraph**, what is Aram’s attitude about his future career? Choices: [(a) He is confident that he will achieve success., (b) He is sure that he will fail, but he wants to try., (c) He is uncertain how things will turn out.] Answer index: 0
@@
Subject: Language Arts Skill: determine-the-main-idea-of-a-passage Description: Read the passage. All over the planet, ancient tombs and temples lie hidden from view. So how do archaeologists know where to begin digging? Thanks to satellites, that question is becoming easier to answer. In the pioneering field of satellite archaeology, satellites four hundred miles above Earth can identify structures buried under soil, covered by modern construction, or obscured by forests. For example, mud bricks used by ancient Egyptians are more dense than soil, and the differences are detected by the satellites. Advanced computer programs map buried settlements and distinguish them from water, vegetation, and farmland. Sarah Parcak, a pioneer in the field, says of the technology, "It allows you to literally strip away vegetation and see entire cities beneath the rain forest canopy. This is the unbelievable future of archaeology." What is the main, or central, idea of the passage? Choices: [(a) Archaeologists must determine where ancient structures are buried before they can begin excavating them., (b) Archaeologists are uncovering ancient hidden structures by using satellite imagery.] Answer index: 1
@@
Subject: Language Arts Skill: use-parallel-structure Description: Complete the sentence. In his novel Nineteen Eighty-Four, George Orwell describes a dystopian London in the hands of a totalitarian government that is paranoid, watchful, and ___. Choices: [(a) mercilessly oppressive, (b) mercilessly oppresses its citizens] Answer index: 0
Subject: Language Arts Skill: use-dictionary-entries-to-determine-correct-usage Description: Look at this dictionary entry. dictionary word: exasperate dictionary pos: verb dictionary inflections: exasperating, exasperated dictionary entry: 1 to make very angry; to irritate Marie was exasperated by Barbara’s constant complaining. dictionary etymology: from Latin (exasperare) Which sentence uses the word **exasperate** correctly? Choices: [(a) Mrs. Hampton realized she was exasperating a bad situation when she failed to pay the parking ticket promptly., (b) Gary’s reluctance to do basic household chores exasperated Khalil and Chase.] Answer index: 1
@@
Subject: Language Arts Skill: trace-an-argument-set-1 Description: Read the text. title: The Connected City The bottom line is that online connectivity is a necessity in today’s high-tech society. Cities across the country are scrambling to keep up with ever-increasing demands for Wi-Fi by the general public. In fact, an Internet Society survey of more than ten thousand Internet users around the world found that eighty-three percent of the respondents agreed that access to the Internet should be considered a basic human right. Whether viewed as a public service or as a way to boost the local economy, people expect advanced technology on the go. Travelers, business owners, and local residents should be able to rely on the local government to help them stay connected. What is the author’s main claim or argument? Choices: [(a) Access to free Wi-Fi throughout a city helps to promote tourism., (b) Free Wi-Fi in cities benefits residents as well as the local economy., (c) Only low-income residents of urban areas should be provided with free Wi-Fi.] Answer index: 1
@@
Subject: Language Arts Skill: which-text-is-most-formal Description: Which expression of thanks is more formal? Choices: [(a) We’re so pleased you could join us on our special day. Thank you again for the wonderful gift., (b) We’re super happy that you could be with us on our special day. Thanks again for the awesome gift!] Answer index: 0
@@
Subject: Language Arts Skill: trace-an-argument-set-2 Description: Read the text. title: Fact or Fiction? Looking beyond the author’s own life events, memoir can inform readers about the world in the same way that other nonfiction can. For example, a memoir written by the head chef of a famous restaurant is likely to educate the reader about the inner workings of a professional kitchen. A memoir composed by a cello player in an esteemed orchestra will probably help the reader learn more about music. And a memoir recalling an adolescence during the 1980s may describe pop culture during that time. Memoir has a way of relaying facts about anything from an occupation to brief fashion trends, all of it meaningful to the author. What is the author’s main claim or argument? Choices: [(a) Memoir is best categorized under the umbrella of literary nonfiction., (b) Memory is too unreliable for memoir to be considered a type of nonfiction., (c) While many memoirs are fact-based, many authors are untrustworthy., (d) The most profitable memoirs are those shown to be the most factual.] Answer index: 0
@@
Subject: Language Arts Skill: choose-the-topic-sentence-that-best-captures-the-main-idea Description: Review the details below. Cell phones can be used to cheat on tests by sending answers to friends or by accessing the Internet. Texting during class is like passing notes and is a waste of time. Choose the best topic sentence to introduce a paragraph containing these details. Choices: [(a) Cell phones in the classroom can have a negative impact on student learning., (b) There are many legitimate arguments for allowing students to have cell phones in class, even if they are disruptive., (c) Cell phones have become a near constant presence at school and at home.] Answer index: 0
@@
Subject: Language Arts Skill: identify-sentence-fragments Description: Is there a sentence fragment? Before signing in, you must first register for an account. Unless you’re already a member of the site. Choices: [(a) yes, (b) no] Answer index: 0
@@
Subject: Language Arts Skill: use-context-as-a-clue-to-the-meanings-of-foreign-expressions Description: What is the meaning of the foreign expression in bold? In 2014, giant moths swarmed **en masse** in parts of Southeast Asia. More than eight hundred sightings were reported in Singapore, while thousands disrupted a soccer match at Darul Makmur Stadium in Malaysia. Choices: [(a) coincidentally, (b) as a group, (c) unexpectedly, (d) fatally] Answer index: 1
@@
Subject: Language Arts Skill: use-the-correct-foreign-expression Description: Complete the text with the correct foreign expression. The Norwood Bistro features an eclectic ___ menu so that diners can select a variety of dishes. Choices: [(a) à la carte, (b) crème de la crème] Answer index: 0
@@
Subject: Language Arts Skill: identify-audience-and-purpose Description: Review the fund-raising pamphlet and then answer the question. Be a hero in an animal’s life! Donate to the Martin County Animal Shelter. Adoption fees cover less than eight percent of our operating expenses. We rely on donations to ensure that every animal receives excellent care. Your generous gift provides food, shelter, and veterinary care to homeless and rescued pets. Please donate today! Which best indicates that the primary audience is "community members who care about animals"? Choices: [(a) The writer indicates that donating is a way for readers to "be a hero in an animal’s life.", (b) The writer argues that the shelter is necessary to keep stray animals off the streets and prevent illnesses., (c) The writer asks readers to donate time as well as money., (d) The writer includes information about funding sources and operating expenses.] Answer index: 0
Subject: Language Arts Skill: classify-figures-of-speech-euphemism-hyperbole-oxymoron-paradox Description: Which figure of speech is used in this text? This movie is so predictable that I’m likely to **die of boredom** if I keep watching it. Choices: [(a) paradox, (b) hyperbole] Answer index: 1
Subject: Language Arts Skill: analyze-short-stories-set-1 Description: Read the story. title: Lake Magic Gramps appeared with dry towels. "I’ve made popcorn and a warm fire. And your mom called. She’s fine." Gramps paused, then added, "Everybody’s fine." What is the main reason Izzie is looking for Champ? Choices: [(a) She hopes Champ will take her brother’s place as a friend to her., (b) She wants to prove to her grandfather that she can find the legendary lake monster., (c) She wants Champ’s "lake magic" to make her family the way it used to be.] Answer index: 2
@@
Subject: Language Arts Skill: explore-words-with-new-or-contested-usages Description: Which text uses the word **nauseous** in its traditional sense? Choices: [(a) Everyone seemed to enjoy the magnolia-scented candle, but it made Walter feel rather **nauseous**., (b) Everyone seemed to enjoy the magnolia-scented candle, but Walter found the smell rather **nauseous**.] Answer index: 1
@@
Subject: Language Arts Skill: analyze-the-development-of-informational-passages-set-2 Description: Read the text. title: Manzanar: A Site of Conscience Adapted from the National Park Service, "Manzanar: A Site of Conscience" What is the main focus of the text? Choices: [(a) how the attack on Pearl Harbor led to war against Japan, (b) the history of the internment of Japanese Americans, with a focus on Manzanar, (c) how Japanese Americans adjusted to everyday life at Manzanar, (d) the history of injustices against Japanese Americans, including internment at Manzanar] Answer index: 1
Subject: Language Arts Skill: analyze-the-development-of-informational-passages-set-1 Description: Read the text. title: Language Change Adapted from the National Science Foundation, "Language and Linguistics: Language Change" What is the main focus of the text? Choices: [(a) why the works of Shakespeare and Chaucer are difficult for modern readers, (b) why different languages evolve at different rates, (c) how and why language changes over time] Answer index: 2
@@
Subject: Language Arts Skill: choose-punctuation-to-avoid-fragments-and-run-ons Description: Which is the best way to complete the text? Zoe traced her family’s lineage back to the eighteenth ___ cross-referenced her grandmother’s stories with a genealogical index at the library. Choices: [(a) century. She, (b) century, furthermore, she, (c) century, she] Answer index: 0
@@
Subject: Language Arts Skill: transition-logically-between-claims-evidence-analysis-and-counterclaims Description: Complete the text with the transition that best connects **the two pieces of supporting evidence**. Parents should limit how much screen time their children are allowed to have each day. A study found that children whose parents limited screen time experienced physical benefits, including better sleep habits and decreased risk of obesity. ___, these children earned better grades in school, were more social and helpful, and were less aggressive with peers. Choices: [(a) In other words, (b) Moreover, (c) For example] Answer index: 1
Subject: Language Arts Skill: choose-the-analysis-that-logically-connects-the-evidence-to-the-claim Description: Read the claim and the supporting evidence. **Claim:** Hunter is remarkably virtuous. **Evidence:** Hunter is generous with his time. Why does the evidence support the claim? Choose the **analysis** that better explains the connection. Choices: [(a) Generosity is a key virtue., (b) Hunter believes generosity need not involve money.] Answer index: 0
@@
Subject: Language Arts Skill: identify-run-on-sentences Description: Is this a run-on sentence? Clinging to a two-point lead in the final minute of the fourth quarter, the Tigers triumphed when their defense recovered a fumble and sealed the victory. Choices: [(a) yes, (b) no] Answer index: 1
Subject: Language Arts Skill: use-words-accurately-and-precisely Description: Complete the text with the better phrase. Karen and Steven are ___ on a documentary film project about the effects of drought on local wildlife. Choices: [(a) collaborating together, (b) working together] Answer index: 1
Subject: Language Arts Skill: classify-figures-of-speech-review Description: Which figure of speech is used in this text? He had hoped to find his missing watch, but **he couldn’t find the time**. Choices: [(a) onomatopoeia, (b) pun] Answer index: 1
Subject: Language Arts Skill: avoid-double-illogical-and-unclear-comparisons Description: Which text best completes the sentence? As a dancer, I studied tap and ballet. In my experience, tap was ___. Choices: [(a) easier to learn than any form of dance, (b) easier to learn, (c) easiest to learn] Answer index: 1
@@
Subject: Language Arts Skill: choose-the-best-evidence-to-support-a-claim Description: Consider this claim: Online colleges prepare students for careers just as well as traditional brick-and-mortar schools do. Which is the strongest evidence to support the claim? Choices: [(a) "I’ve hired two people recently who received their college degrees through online programs, and the quality of their performance is as good as, if not better than, that of their peers," reported business owner Fred Browning., (b) Online college programs often provide career preparation services that are equivalent to those provided by traditional college programs, including interview preparation, job fairs, and even internship placement., (c) Many students choose to gain some credits through an online college program and then successfully transfer to a brick-and-mortar institution to complete a degree.] Answer index: 1
@@
Subject: Language Arts Skill: analyze-short-stories-set-2 Description: Read the story. title: Water Rising Polara looked to the horizon, at the unmistakable sparkle of the approaching sea. She wasn’t sure how brave she felt, but she was willing. "I am," she answered. Based on the **first and second paragraphs**, what internal conflict does Polara struggle with? Choices: [(a) the duty to stand watch versus her inclination to daydream, (b) the pull of the unfamiliar versus her desire to stay at Shublik, (c) respect for Commander Axel versus her irritation at his order, (d) the enjoyment of the view versus her fear about the water on the horizon] Answer index: 1
@@
Subject: Language Arts Skill: use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing Description: Which statement primarily appeals to **logos**, or reason, to support the position that **more money does not always lead to increased happiness**? Choices: [(a) Like most people, I enjoy the feeling of buying a new gadget, but it turns out that the happiness that we get from a new purchase doesn’t last long., (b) While it’s true that a lack of money can cause stress and unhappiness, it does not necessarily follow that money will lead to more happiness after basic needs are met.] Answer index: 1
Subject: Language Arts Skill: choose-the-word-whose-connotation-and-denotation-best-match-the-sentence Description: Complete the sentence with the word that best fits the overall meaning and tone. Dan appreciated the gallery owner’s ___ efforts to include artwork from the local community. Choices: [(a) calculating, (b) deliberate] Answer index: 1
@@
Subject: Language Arts Skill: evaluate-counterclaims Description: Consider this claim: Veggie burgers are every bit as tasty and satisfying as regular beef burgers. Now consider how someone might argue against this. Which sentence presents the stronger and more reasonable counterclaim to the above claim? Choices: [(a) Veggie burgers have no taste whatsoever and cannot compare in any way to a delicious beef burger., (b) Regular beef burgers have more good-tasting fat and filling protein than veggie burgers.] Answer index: 1